OpenStudy (anonymous):

What is the slope of the line 2x – 5y = -10?

OpenStudy (anonymous):

2x+10=5y (2/5)x+2=y y=(2/5)x+2 This is in the form of y=mx+c m=slope From the two equations, m=2/5